# Sequential Change Detection under Markov Setup with Unknown Prechange and Postchange Distributions

### v2

@@ Abstract
In this work we extend the results developed in 2022 for a sequential change detection algorithm making use of Page’s CUSUM statistic,empirical distribution as an estimate of pre-change distribution and a universal code as a tool for estimating the post change distribution, from IID to Markov setup.
@@ I Introduction
Change detection has numerous applications including but not limited to industrial quality control,fault detection in computer communication networks and regime shift detection in financial markets.
@@
Page introduced the Cumulative Sum Test (CUSUM) to address the problem of sequential change detection in 1954 [4].Lorden proved the asymptotic optimality of CUSUM for memoryless sources in minimax framework [6] while Moustadikes proved its exact optimality [15]. Shiryaev [5, 10] introduced a different approach to the change detection problem under Bayesian framework which has been extensively studied in literature.Lorden’s result was extended for sources with memory by Bansal and Kazakos [16] who assume independence of observations before and after change which was further weakened in Lai’s extension [20].[8, 21] are other extensions of CUSUM beyond i.i.d. framework.
@@
Malik and Bansal [26] considered the case where both pre-change and post-change distributions are unknown for memoryless sources.They further modified the test introduced by Jacob and Bansal [24] ,by using an empirical estimate in the place of true prechange distribution without affecting the post-change distribution estimate and proved the asymptotic optimality of the proposed test.
In this paper, we extend the modification demonstrated by Malik and Bansal [26] for the more general class of stationary and ergodic finite-order Markov sources that satisfy Berk’s mixing conditions[7] and make use of Bansal and Kazakos’ extension [16] to study the performance of proposed test and establish its optimality in the asymptotic regime where the rate of false alarm goes to 0.
Our main contributions in the paper are:
1.Upper bound on probability of error for proposed test in Lemma 3.
2.Showed that the proposed test will terminate definitely under alternate hypothesis in Lemma 4.
3.Evaluated the performance of proposed test by studying linear cost function of the number of samples required to achieve a specific probability of error in Theorem 2 and the linear cost function of detection delay required to achieve a specific false alarm rate in Theorem 5.
4.Established the optimality of proposed test in the asymptotic regime in Theorem 6.
Section II introduces the preliminaries required for the rest of the paper.In Section III, we introduce the empirical estimate,the modified test and study behaviour of proposed test under pre-change and post-change distribution.In Section IV, we study the performance of modified test and prove its asymptotic optimality.
@@ II Preliminaries
Let $X=\{X_{n}\}_{n=1}^{\infty}$ be a source with finite alphabet $\mathcal{X}$.A sequence of source symbols $x_{i},x_{i+1},\dots,x_{j}$ is denoted by $x_{i}^{j}$ and corresponding random variables $X_{i},X_{i+1},\dots,X_{j}$ is denoted by $X_{i}^{j}$.All the logarithms will be taken to the base 2 in the remaining, unless specified otherwise.
A fixed-to-variable code (FV code) is a map $\phi_{n}:X^{n}\to\{0,1\}^{*}$, where $\{0,1\}^{*}$ denotes the set of all finite-length strings.Let $L(x_{1}^{n})=|\phi_{n}(x_{1}^{n})|$ denotes the length of the codeword for a given sequence $x_{1}^{n}$ which satisfies prefix condition.Therefore, Kraft inequality holds, given as
$\sum_{x_{1}^{n}\in X^{n}}2^{-L(x_{1}^{n})}\leq 1$
Let $\mu$ be the distribution of $X=\{X_{n}\}_{n=1}^{\infty}$.The entropy rate for stationary $\mu$ is
$H(\mu)=-\lim_{n\to\infty}\frac{1}{n}\sum_{x_{1}^{n}\in X^{n}}\mu(x_{1}^{n})\log\mu(x_{1}^{n})$
For a stationary ergodic source, Shannon-Mcmillian-Brieman theorem states
$\lim_{n\to\infty}\frac{1}{n}\log\mu(x_{1}^{n})=H(\mu)\quad\mu\text{ a.s.}$
A FV code with length function $L(x_{1}^{n})$ is said to be pointwise universal if
@@
$R_{L}(n)=\sup_{\mu\in\mathcal{M}}\max_{x_{1}^{n}\in X^{n}}\left(L(x_{1}^{n})+\log\mu(x_{1}^{n})\right)\sim o(n)$
The condition for pointwise universality (2.4) is equivalent to having $\mu(J_{\epsilon}<\infty)=1$ for all $\epsilon>0$, where
$J_{\epsilon}=\sup_{n\geq 1}\left\{n:\left|\frac{1}{n}\left(L(x_{1}^{n})+\log\mu(x_{1}^{n})\right)\right|\geq\epsilon\right\}$ (2.6)
A stronger version of the convergence condition (2.4) requires that
$\mathbb{E}_{\mu}(J_{\epsilon})<\infty\quad\text{for all}\quad\epsilon>0$ (2.7)
Strongly pointwise universal (2.5) codes satisfy this condition which is called 1-quick universal by Jacob [25], for every source in the class $\mathcal{M}$.For Markov sources upto a fixed order, the existence of codes that satisfy all these conditions are known(Lempel-Ziv codes (LZ-77, LZ-78) [9, 11] and the Context Tree Weighting code) [19] and we consider such codes only in our paper.
Theorem 1(Barron)[13, 18]: Let $\mu$ be a stationary ergodic process and $\nu$ be a stationary Markov process of finite order such that $\mu\ll\nu$.Then
@@
where $D(\mu||\nu)$ is the divergence rate.
The pointwise convergence in the above theorem is equivalent to having $\mu(K_{\epsilon}<\infty)=1\quad\text{for all}\quad\epsilon>0$ where
$K_{\epsilon}=\sup_{n\geq 1}\left\{n:\left|\frac{1}{n}\log\frac{\mu(x_{1}^{n})}{\nu(x_{1}^{n})}-D(\mu||\nu)\right|\geq\epsilon\right\}\quad$
A stronger version of this convergence, called 1-quick convergence of log-likelihood ratios according to Lai’s terminology [12], requires that
@@
For a fixed sample hypothesis test with $H_{0}:\mu=\mu_{0}$ as the null hypothesis and $H_{1}:\mu=\mu_{1}$ as the alternative, consider the discriminant function
$h(x_{1}^{n})=\log\mu_{1}(x_{1}^{n})-\log\mu_{0}(x_{1}^{n})-n\lambda$
After observing $n$ samples $x_{1}^{n}$, we reject the null hypothesis if $h(x_{1}^{n})\geq 0$. The following lemma characterizes the performance of this test.
Lemma 1 [3, 17]:If $\mu_{0}$ and $\mu_{1}$ are stationary Markov process of finite order and stationary ergodic process such that $\mu_{1}\ll\mu_{0}$ and $D(\mu_{1}\|\mu_{0})>\lambda$. For the test induced by the discriminant function $h(x_{1}^{n})$ we have:
$\mu_{0}\big(h(x_{1}^{n})\geq 0\big)$  $\;\leq\;2^{-n\lambda}$ (2.12)
$\lim_{n\to\infty}\mu_{1}\big(h(x_{1}^{n})\geq 0\big)$ $=1.$ (2.13)
Equation (2.13) remains true even if the threshold 0 is replaced by any other constant.The lemma follows immediately from the Shannon–McMillan–Breiman theorem.
@@
$q_{n}(\Delta)=\mu_{0}\left(\frac{1}{n}\log\frac{\mu_{0}(x_{1}^{n})}{\mu_{1}(x_{1}^{n})}<\Delta\right)$
For $\Delta<D(\mu_{1}\|\mu_{0})$, let
$\lim_{n\to\infty}np_{n}(\Delta)=0\quad\text{and}\quad\sum_{n\geq 1}p_{n}(\Delta)<\infty$
For $\Delta<D(\mu_{0}\|\mu_{1})$, let
$\lim_{n\to\infty}nq_{n}(\Delta)=0\quad\text{and}\quad\sum_{n\geq 1}q_{n}(\Delta)<\infty$
@@ II-A Previous Work
For the case where both pre-change $\mu_{0}$ and post-change $\mu_{1}$ distributions are known, Page proposed a test [4] which is given as follows.
@@
$\max_{k\leq n}\left(\log\mu_{1}(x_{k}^{n})-\log\mu_{0}(x_{k}^{n})\right)\geq\log\gamma,$ (2.18)
@@
Jacob–Bansal Page CUSUM Test (Test 2)[24, 25]: A modification of Page’s CUSUM test was proposed for cases where the post-change distribution $\mu_{1}$ is unknown for i.i.d. sources and later extended for sources with memory satisfying (2.10),(2.16),(2.17). The modified CUSUM Test, where a strongly pointwise universal code is used to estimate the post-change distribution, is given as follows.
@@
$\max_{k\leq n}\left(-L(x_{k}^{n})-\log\mu_{0}(x_{k}^{n})-n\lambda\right)\geq\log\gamma,$ (2.19)
@@
Here, $\lambda$ was introduced by the authors in order to ensure a negative drift in the test statistic in the pre-change phase, as is the case in the classical Page’s test in the parametric setup.
@@
$N_{0}(\alpha)=\inf\left\{n:-L(x_{1}^{n})-\log\mu_{0}(x_{1}^{n})-n\lambda\geq-\log\alpha\right\}.$ (2.20)
@@
Test($\gamma>1$$\lambda>0$):Start with $n=1$,for samples $x_{n_{0}+1},x_{n_{0}+2},\dots,x_{n_{0}+n}$.If:
$\max_{1\leq k\leq n}\left(-L(x_{n_{0}+k}^{n_{0}+n})-\log\hat{\mu}_{0}({x}_{n_{0}+k}^{n_{0}+n})-n\lambda\right)\geq\log\gamma,$ (2.21)
@@ III Modified CUSUM Test For Markov
We consider stationary and ergodic mutually independent first order Markov sources $\mu_{0}$ and $\mu_{1}$ satisfying (2.10),(2.16),(2.17) and $\mu_{1}\ll\mu_{0}$ as well as codes satisfying (2.7) for analysis which can be generalised for any finite order Markov sources satisfying these conditions.We also assume that if $\mu_{0}\in\mathcal{M}_{0}$ and $\mu_{1}\in\mathcal{M}_{1}$ then $\min_{\begin{subarray}{c}\mu_{0}\in\mathcal{M}_{0}\\ \mu_{1}\in\mathcal{M}_{1}\end{subarray}}D(\mu_{1}\|\mu_{0})=\lambda_{0}>0$.
@@ III-A Estimating Pre-change Distribution
For a stream of source symbols with no change up to $n_{0}$, the test has two stages. In the first stage, we estimate the prechange distribution using the first $n_{0}$ symbols, written as $x_{1},x_{2},\dots,x_{n_{0}}$ and the subsequent symbols as $x_{n_{0}+1},x_{n_{0}+2},\dots,x_{n_{0}+n}$. The prechange distribution estimate, $\hat{\mu}_{0}$, is calculated from initial sample of $n_{0}$ symbols as,
$\hat{\mu}_{0}(b,a)=\frac{N(b,a)\mid x_{1}^{n_{0}}}{n_{0}-1}$
$\hat{\mu}_{0}(b)=\frac{N(b)\mid x_{1}^{n_{0}}}{n_{0}}$
@@
Here, $N(b)\mid x_{1}^{n_{0}}$ represents the number of times the symbol $b$ appears in the sequence $x_{1}^{n_{0}}$, whereas $N(b,a)\mid x_{1}^{n_{0}}$ represents the number of times the symbols $b,a$ jointly appear in the sequence $x_{1}^{n_{0}}$.Also, $\hat{\mu}_{0}(b)$ is the fraction of times $b$ appears in the sequence $x_{1}^{n_{0}}$, whereas $\hat{\mu}_{0}(b,a)$ is the fraction of times $(b,a)$ appear in the sequence $x_{1}^{n_{0}}$ jointly.
@@
Since ${\mu}_{0}$ is a stationary and ergodic finite-order Markov source, by Ergodic Theorem, we can say that there exists an $n_{0}$ such that,
@@ III-B Test
Using the estimate $\hat{\mu}_{0}$ and extending the Modified CUSUM Test for Markov sources by defining it as,
Test($\gamma>1$$\lambda>0$): Start with $n=1$ for samples $x_{n_{0}+1},x_{n_{0}+2},\dots,x_{n_{0}+n}$.If:
$\max_{1\leq k\leq n}\left(-L(x_{n_{0}+k}^{n_{0}+n})-\log\hat{\mu}_{0}({x}_{n_{0}+k}^{n_{0}+n})-n\lambda\right)\geq\log\gamma$
Stop the test and decide that a change has occurred. Otherwise, sample for $x_{n_{0}+n+1}$ and continue the test. The upward drift rate of this test is $D(\mu_{1}||\hat{\mu}_{0})-\lambda$ and the downward drift rate of this test is $\lambda-D(\mu_{0}||\hat{\mu}_{0})$. Hence, for this test to work, need:
@@
$N(\alpha)=\inf\left\{n:\left(-L(x_{n_{0}+k}^{n_{0}+n})-\log\hat{\mu}_{0}(x_{n_{0}+k}^{n_{0}+n})-n\lambda\right)\geq-\log\alpha\right\}$ (3.10)
III-C Behaviour Under Prechange Distribution
Introducing the following notation to study the expected behaviour of the auxillary stopping time $N(\alpha)$ under prechange distribution $\mu_{0}$ and postchange distribution $\mu_{1}$ with respect to threshold parameter $\alpha$.
$f_{n}(x_{n_{0}+1}^{n_{0}+n})=\frac{\log\frac{\mu_{0}(x_{n_{0}+1}^{n_{0}+n})}{\hat{\mu}_{0}(x_{n_{0}+1}^{n_{0}+n})}}{n}$
$l_{n}(x_{1}^{n_{0}+n})=-L(x_{n_{0}+1}^{n_{0}+n})-\log\hat{\mu}_{0}(x_{n_{0}+1}^{n_{0}+n})-n\lambda$
$B_{n}\coloneq\{x_{1}^{n_{0}+n}:l_{k}(x_{1}^{n_{0}+k})<-\log\alpha,\;1\leq k<n,\;l_{n}(x_{1}^{n_{0}+n})\geq-\log\alpha\}$
$B=\bigcup_{n=1}^{\infty}B_{n}$
$E_{n}\coloneq\left\{x_{1}^{n_{0}+n}:\begin{array}[]{l}\big|\hat{\mu}_{0}(a\mid b)-\mu_{0}(a\mid b)\big|<\delta\quad\forall\,(a,b)\in\mathcal{X}^{2},\\[6.0pt] \big|\hat{\mu}_{0}(b)-\mu_{0}(b)\big|<\delta^{\prime}\quad\forall\,b\in\mathcal{X}\end{array}\right\},$
$\text{where }\hat{\mu}_{0}\text{ is based on samples }x_{1}^{n_{0}}.$
$E=\bigcup_{n=1}^{\infty}E_{n}$
Lemma 2: If $x_{1}^{n_{0}+n}\in E_{n}$, then $f_{n}(x_{n_{0}+1}^{n_{0}+n})$ is bounded above for all $n$ as $f_{n}\leq W_{n}$ where $\beta_{1}=\frac{\mu_{0}(a^{\prime}\mid b^{\prime})}{\mu_{0}(a^{\prime}\mid b^{\prime})-\delta},\beta_{2}=\frac{\mu_{0}(b^{\prime})}{\mu_{0}(b^{\prime})-\delta^{\prime}}$ and $(b^{\prime},a^{\prime})$ is such that $(b^{\prime},a^{\prime})$ is the minimum transition probability symbol pair and $b^{\prime}$ is such that $b^{\prime}$ is the minimum probability symbol.As $\delta\to 0$$\delta^{\prime}\to 0$ and $n_{0}\to\infty$, we have $\beta_{1}\to 1$ and $\beta_{2}\to 1$ .
Proof:As $x_{i}$’s follow first order markov process satisfying (2.10),(2.16),(2.17), can write $f_{n}\big(x_{n_{0}+1}^{n_{0}+n}\big)$ as,

### v3

# Sequential Change Detection under Markov Setup with Unknown Pre-change and Post-change Distributions

@@ Abstract
In this work we extend the results developed in 2022 for a sequential change detection algorithm making use of Page’s CUSUM statistic, empirical distribution as an estimate of pre-change distribution and a universal code as a tool for estimating the post-change distribution, from i.i.d. to Markov setup.
@@ I Introduction
Change detection has numerous applications including industrial quality control, and fault detection in computer communication networks and has been extensively studied in literature. Sequential hypothesis testing and sequential change detection have a long history.
@@
Page introduced the Cumulative Sum Test (CUSUM) to address the problem of sequential change detection in 1954 [4]. Lorden proved the asymptotic optimality of CUSUM for memoryless sources in minimax framework [6] while Moustakides proved its exact optimality [15]. Shiryaev [5, 10] and Pollak [14] proposed distinct approaches for the change detection problem under Bayesian and non-Bayesian framework respectively. Lorden’s result was extended for sources with memory by Bansal and Kazakos [16] who assume independence of observations before and after change which was further weakened in Lai’s extension [20]. [8, 21] are other works addressing the unknown post-change parameter case.
@@
Malik and Bansal [26] considered the case where both pre-change and post-change distributions are unknown for memoryless sources. They further modified the test introduced by Jacob and Bansal [24], by using an empirical estimate in the place of true pre-change distribution without affecting the post-change distribution estimate and proved the asymptotic optimality of the proposed test.
In this paper, we extend the modification demonstrated by Malik and Bansal [26] for the more general class of stationary and ergodic finite-order Markov sources that satisfy Berk’s mixing conditions[7] and make use of Bansal and Kazakos’ extension [16] to study the performance of proposed test and establish its optimality in the asymptotic regime where the rate of false alarm goes to 0. We evaluate the performance of the proposed test by providing an upper bound on its probability of error, showing that it will terminate with probability one under the alternative hypothesis, and studying a cost function for the number of samples required to achieve a specific probability of error as well as a cost function for the detection delay required to achieve a specific false alarm rate.
Section II introduces the preliminaries required for the rest of the paper. In Section III, we introduce the empirical estimate, the modified test and study the behaviour of proposed test under pre-change and post-change distribution. In Section IV, we study the performance of modified test and prove its asymptotic optimality and conclude the paper in Section V.
@@ II Preliminaries
Let $X=\{X_{n}\}_{n=1}^{\infty}$ be a random process (source) taking values from a finite set called the source alphabet $\mathcal{X}$. A sequence of source symbols $x_{i},x_{i+1},\dots,x_{j}$ is denoted by $x_{i}^{j}$ and corresponding random variables $X_{i},X_{i+1},\dots,X_{j}$ is denoted by $X_{i}^{j}$. All the logarithms will be taken to the base 2 in the remaining, unless specified otherwise.
For a fixed n, a fixed-to-variable code (FV code) is a one-to-one mapping $\phi_{n}:X^{n}\to\{0,1\}^{*}$, where $\{0,1\}^{*}$ denotes the set of all finite-length strings. Let $L(x_{1}^{n})=|\phi_{n}(x_{1}^{n})|$ denote the length of the codeword for a given sequence $x_{1}^{n}$ which satisfies prefix condition. Therefore, Kraft inequality holds, given as,
$\sum_{x_{1}^{n}\in X^{n}}2^{-L(x_{1}^{n})}\leq 1.$
Let $\mu$ be the distribution of $X=\{X_{n}\}_{n=1}^{\infty}$. The entropy rate for stationary $\mu$ is well defined [23]. For a stationary ergodic process, entropy rate of the source is a well-known lower bound for the performance of any source coding algorithm [23]. Optimal codes reach this lower bound. A FV code with length function $L(x_{1}^{n})$ is said to be pointwise universal if,
@@
$R_{L}^{n}=\sup_{\mu\in\mathcal{M}}\max_{x_{1}^{n}\in X^{n}}\left(L(x_{1}^{n})+\log\mu(x_{1}^{n})\right)\sim o(n).$
The condition for pointwise universality (2) is equivalent to having $\mu(J_{\epsilon}<\infty)=1$ for all $\epsilon>0$, where,
$J_{\epsilon}=\sup_{n\geq 1}\left\{n:\left|\frac{1}{n}\left(L(x_{1}^{n})+\log\mu(x_{1}^{n})\right)\right|\geq\epsilon\right\}.$ (4)
A stronger version of the convergence in condition (2) requires that,
$\mathbb{E}_{\mu}(J_{\epsilon})<\infty\quad\text{for all}\quad\epsilon>0.$ (5)
Strongly pointwise universal codes (3) satisfy this condition which is called 1-quick universal by Jacob [25], for every source in the class $\mathcal{M}$. For Markov sources upto a fixed order, the existence of codes that satisfy conditions (3) and (5) are known (Lempel-Ziv codes (LZ-77, LZ-78) [9, 11] and the Context Tree Weighting code [19] ) and we consider such codes only in our paper. The divergence rate for two stationary processes $\mu$ and $\nu$ is defined as,
$D(\mu||\nu)=\lim_{n\to\infty}\frac{1}{n}\sum_{x_{1}^{n}\in X^{n}}\mu(x_{1}^{n})\log\frac{\mu(x_{1}^{n})}{\nu(x_{1}^{n})}$
provided the limit exists.
Theorem 1 (Barron)[13, 18]: Let $\mu$ be a stationary, ergodic process and $\nu$ be a stationary, finite-order Markov process such that $\mu\ll\nu$. Then,
@@
where $D(\mu||\nu)$, the divergence rate is well defined.
The pointwise convergence in the above theorem is equivalent to having $\mu(K_{\epsilon}<\infty)=1\quad\text{for all}\quad\epsilon>0$ where,
$K_{\epsilon}=\sup_{n\geq 1}\left\{n:\left|\frac{1}{n}\log\frac{\mu(x_{1}^{n})}{\nu(x_{1}^{n})}-D(\mu||\nu)\right|\geq\epsilon\right\}\quad.$
A stronger version of this convergence, called 1-quick convergence of log-likelihood ratios according to Lai’s terminology [12, 25], requires that,
@@
For a fixed sample hypothesis test with the null hypothesis $H_{0}:\mu=\mu_{0}$ and the alternative $H_{1}:\mu=\mu_{1}$, consider the discriminant function,
$h(x_{1}^{n})=\log\mu_{1}(x_{1}^{n})-\log\mu_{0}(x_{1}^{n})-n\lambda.$
After observing $n$ samples $x_{1}^{n}$ we reject the null hypothesis if $h(x_{1}^{n})\geq 0$. The performance of this test is characterized by the following lemma.
Lemma 1 [3, 17]: If $\mu_{0}$ and $\mu_{1}$ are stationary, finite-order Markov process and stationary, ergodic process such that $\mu_{1}\ll\mu_{0}$ and $D(\mu_{1}\|\mu_{0})>\lambda$. For the test induced by the discriminant function $h(x_{1}^{n})$ we have:
$\mu_{0}\big(h(x_{1}^{n})\geq 0\big)$  $\;\leq\;2^{-n\lambda}$ (11)
$\lim_{n\to\infty}\mu_{1}\big(h(x_{1}^{n})\geq 0\big)$ $=1.$ (12)
Equation (12) remains true even if the threshold 0 is replaced by any other constant.
@@
$q_{n}(\Delta)=\mu_{0}\left(\frac{1}{n}\log\frac{\mu_{0}(x_{1}^{n})}{\mu_{1}(x_{1}^{n})}<\Delta\right).$
For $\Delta<D(\mu_{1}\|\mu_{0})$,
$\lim_{n\to\infty}np_{n}(\Delta)=0\quad\text{and}\quad\sum_{n\geq 1}p_{n}(\Delta)<\infty.$
For $\Delta<D(\mu_{0}\|\mu_{1})$,
$\lim_{n\to\infty}nq_{n}(\Delta)=0\quad\text{and}\quad\sum_{n\geq 1}q_{n}(\Delta)<\infty.$
@@ II-A Previous Work
For the case where both pre-change distribution $\mu_{0}$ and post-change $\mu_{1}$ distribution are known, Page proposed a test [4] which is given as follows.
@@
$\max_{k\leq n}\left(\log\mu_{1}(x_{k}^{n})-\log\mu_{0}(x_{k}^{n})\right)\geq\log\gamma,$ (17)
@@
Jacob–Bansal Page CUSUM Test (Test 2)[24, 25]: A modification of Page’s CUSUM test was proposed for cases where the post-change distribution $\mu_{1}$ is unknown for i.i.d. sources and later extended for sources with memory satisfying (9), (15), (16). The modified CUSUM Test, where a strongly pointwise universal code is used to estimate the post-change distribution, is given as follows.
@@
$\max_{k\leq n}\left(-L(x_{k}^{n})-\log\mu_{0}(x_{k}^{n})-n\lambda\right)\geq\log\gamma,$ (18)
@@
Authors introduced $\lambda$ here to ensure a negative drift in the test statistic in the pre-change regime, as is the case in the classical Page’s test in the parametric setup.
@@
$N_{0}(\alpha)=\inf\left\{n:-L(x_{1}^{n})-\log\mu_{0}(x_{1}^{n})-n\lambda\geq-\log\alpha\right\}.$ (19)
@@
Test ($\gamma>1$, $\lambda>0$) : Start with $n=1$, for samples $x_{n_{0}+1},x_{n_{0}+2},\dots,x_{n_{0}+n}$. If:
$\max_{1\leq k\leq n}\left(-L(x_{n_{0}+k}^{n_{0}+n})-\log\hat{\mu}_{0}({x}_{n_{0}+k}^{n_{0}+n})-n\lambda\right)\geq\log\gamma,$ (20)
@@ III Modified CUSUM Test For Markov
We consider stationary, ergodic, mutually independent, first-order Markov sources $\mu_{0}$ and $\mu_{1}$ satisfying (9), (15), (16) and $\mu_{1}\ll\mu_{0}$ as well as universal codes satisfying (5) for analysis. We also assume that if $\mu_{0}\in\mathcal{M}_{0}$ and $\mu_{1}\in\mathcal{M}_{1}$ then $\min_{\begin{subarray}{c}\mu_{0}\in\mathcal{M}_{0}\\ \mu_{1}\in\mathcal{M}_{1}\end{subarray}}D(\mu_{1}\|\mu_{0})=\lambda_{0}>0$.
@@ III-A Estimating Pre-change Distribution
For a stream of source symbols with no change up to $n_{0}$, the test has two stages. In the first stage, we estimate the pre-change distribution using the first $n_{0}$ symbols, written as $x_{1},x_{2},\dots,x_{n_{0}}$ and the subsequent symbols as $x_{n_{0}+1},x_{n_{0}+2},\dots,x_{n_{0}+n}$. The prechange distribution estimate, $\hat{\mu}_{0}$, is calculated from initial sample of $n_{0}$ symbols as,
$\hat{\mu}_{0}(b,a)=\frac{N(b,a)\mid x_{1}^{n_{0}}}{n_{0}-1}\quad\forall(b,a)\in\mathcal{X}^{2}$
$\hat{\mu}_{0}(b)=\frac{N(b)\mid x_{1}^{n_{0}}}{n_{0}}\quad\forall b\in\mathcal{X}$
@@
[…]
@@
Since ${\mu}_{0}$ is a stationary and ergodic first-order Markov source, by Ergodic Theorem, we can say that there exists an $n_{0}$ such that,
@@ III-B Test
Using the estimate $\hat{\mu}_{0}$, we extend the modified CUSUM Test for Markov sources by defining it as,
Test ($\gamma>1$): Start with $n=1$ for samples $x_{n_{0}+1},x_{n_{0}+2},\dots,x_{n_{0}+n}$. If
$\max_{1\leq k\leq n}\left(-L(x_{n_{0}+k}^{n_{0}+n})-\log\hat{\mu}_{0}({x}_{n_{0}+k}^{n_{0}+n})-n\lambda\right)\geq\log\gamma,$
stop the test and decide that a change has occurred. Else, sample for $x_{n_{0}+n+1}$ and continue the test. The upward drift rate and downward drift rate of this test are $D(\mu_{1}||\hat{\mu}_{0})-\lambda$ and $\lambda-D(\mu_{0}||\hat{\mu}_{0})$ respectively. Hence, for this test to work, we need:
@@
$N(\alpha)=\inf\Bigl\{n:\Bigl($ $-L(x_{n_{0}+1}^{n_{0}+n})-\log\hat{\mu}_{0}(x_{n_{0}+1}^{n_{0}+n})$ (30)
$-n\lambda\Bigr)\geq-\log\alpha\Bigr\}.$
III-C Behaviour Under Pre-change Distribution
Introducing the following notation to study the expected behaviour of the auxillary stopping time $N(\alpha)$ under pre-change distribution $\mu_{0}$ and post-change distribution $\mu_{1}$.
$f_{n}(x_{n_{0}+1}^{n_{0}+n})\triangleq\frac{\log\frac{\mu_{0}(x_{n_{0}+1}^{n_{0}+n})}{\hat{\mu}_{0}(x_{n_{0}+1}^{n_{0}+n})}}{n}$
$l_{n}(x_{1}^{n_{0}+n})\triangleq-L(x_{n_{0}+1}^{n_{0}+n})-\log\hat{\mu}_{0}(x_{n_{0}+1}^{n_{0}+n})-n\lambda$
$B_{n}\triangleq\{x_{1}^{n_{0}+n}:l_{k}(x_{1}^{n_{0}+k})<-\log\alpha,1\leq k<n,\;l_{n}(x_{1}^{n_{0}+n})$
$\geq-\log\alpha\}$
$B\triangleq\bigcup_{n=1}^{\infty}B_{n}$
$E_{n}\triangleq\Bigl\{x_{1}^{n_{0}+n}:$ $\lvert\hat{\mu}_{0}(a\mid b)-\mu_{0}(a\mid b)\big|<\delta\quad\forall\,(a,b)\in\mathcal{X}^{2},$
$\lvert\hat{\mu}_{0}(b)-\mu_{0}(b)\big|<\delta^{\prime}\quad\forall\,b\in\mathcal{X}\Bigr\}$
 $\text{ where }\hat{\mu}_{0}\text{ is based on samples }x_{1}^{n_{0}}$
$E\triangleq\bigcup_{n=1}^{\infty}E_{n}$
Lemma 2: If $x_{1}^{n_{0}+n}\in E_{n}$, then $f_{n}(x_{n_{0}+1}^{n_{0}+n})$ is bounded above for all $n$ as $f_{n}\leq W_{n}=\frac{(n-1)\log(\beta_{1})+\log(\beta_{2})}{n}$, where $\beta_{1}=\frac{\mu_{0}(a^{\prime}\mid b^{\prime})}{\mu_{0}(a^{\prime}\mid b^{\prime})-\delta},\beta_{2}=\frac{\mu_{0}(b^{\prime})}{\mu_{0}(b^{\prime})-\delta^{\prime}}$ and $(b^{\prime},a^{\prime})$ is such that $(b^{\prime},a^{\prime})$ is the minimum transition probability symbol pair and $b^{\prime}$ is such that $b^{\prime}$ is the minimum probability symbol. As $\delta\to 0$, $\delta^{\prime}\to 0$ and $n_{0}\to\infty$, we have $\beta_{1}\to 1$ and $\beta_{2}\to 1$.
Proof: As $x_{i}$’s follow first order Markov process, we can write $f_{n}\big(x_{n_{0}+1}^{n_{0}+n}\big)$ as,